Exercise 10-8 Allocating product costs between ending inventory and cost of goods sold LO 10-3 Vernon Manufacturing Company ba. Total product cost b. Total cost of ending inventory c. Total cost of goods sold
Answer #1

Total product cost

= 3200+3530+5388

= 12,118

Ending inventory = Units produced - Units sold = 1660-1150 = 510

= (12118/1660)*510

= 3723

Cost of Goods Sold

= (12118/1660)*1150

= 8395